4. Adam’s Inc.’s outstanding common stock is currently selling in the market for $28. Dividends of $1.80 per share were paid last year, and the company expects annual growth of 6%. (a) What is the value of the stock to you, given a 11% required rate of return? (b) Determine the expected rate of return for the stock. (c) Should you purchase this stock? Why?
Answer #1

a.

As per CAPM Model,

Value of Stock = D0(1 + g)/(r - g)

So,

Value of Stock = 1.80(1.06)/(0.11 - 0.06)

Value of Stock = $38.16

b.

Calculating Expected Return,

Stock Price = D0(1 + g)/(r - g)

r = D0(1 + g)/Price + g

r = 1.80(1.06)/28 + 0.06

r = 12.81%

c.

One should purchase the stock, because,

Expected Return(12.81%) > Required Return(11%)